Write the given expression in the form f(x)=a(x-h)^2+k. Identify the vertex
f(x)=1/3x^2+x+1
If you could, please show me step by step. Thanks!

Answer by ewatrrr(24785)


Hi
Write the given expression in the standard vertex form f(x)=a(x-h)^2+k
where(h,k) is the vertex
f(x)=1/3x^2+x+1 |completing the square to put into vertex form
f(x)=1/3[x^2+3x)+1
f(x)=1/3[x +3/2)^2 -9/4]+1
f(x)=1/3(x +3/2)^2 -3/4 + 1
f(x)=1/3(x +3/2)^2 + 1/4 |vertex is Pt(-3/2,1/4)